OpenStudy (anonymous):

Hiro and lisa are photographers .They charge a standard fee for taking pictures. Theres also a charge for each set of prints ordered.So the number of sets of prints determines the total cost. sets of prints(s) Total Cost 0 $30 1 $31 2 $32 3 $33 Let s stand for the sets of prints ordered. Write the rule that describes a students total cost.

OpenStudy (anonymous):

Please help!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (snowsurf):

Do you need an equation or just describing?

OpenStudy (anonymous):

equation

OpenStudy (anonymous):

@snowsurf\

OpenStudy (anonymous):

Hello ? @snowsurf

OpenStudy (snowsurf):

Okay. Lets have S be sets of prints and let T be total cost. So we will have an equation like this T = mS + b. Its like the linear equation y = mx + b. Now lets you two of the points you have. (0, 30) (1, 31) 30 = m(0) + b ----> b = 30 31 = m(1) + b -----> m = 31 - b Plug in b into the other equation. So you will have m = 31 - 30 ----> m =1 So we have b=30 and m =1 T = S + 30 will be your equation.